# Toward Universal Laws of Outlier Propagation

### v2

@@ 2.4 Algorithmic Markov condition and independence of mechanisms
Let $x_{1},\dots,x_{n}$ be binary words describing objects whose causal relations are given by the DAG $G$. Then the joint complexity of $(x_{1},\dots,x_{n})$ decomposes as
@@
Furthermore, for any three sets $R,S,T$ of nodes, we have
$I(R:S\mid T^{*})\stackrel{{\scriptstyle+}}{{=}}0,$
whenever $R$ and $S$ are d-separated by $T$.
Janzing and Schölkopf [2010] argue that within a causal Bayesian network, where nodes represent random variables, the mechanisms $P_{X_{j}|\operatorname{{PA}}_{j}}$ constitute independent information-bearing objects. Given $m$ observations of $n$-tuples $(x^{1}_{1},\dots,x_{n}^{1}),\dots,(x^{m}_{1},\dots,x_{n}^{m})$, they define a DAG $G^{m}$ that connects the $n$ mechanisms $P_{X_{i}|\operatorname{{PA}}_{i}}$ with the $n\times m$ observations in the following way: Each $x_{i}^{l}$ has the observations $\operatorname{{pa}}^{l}_{i}$ and the node $P_{X_{i}|\operatorname{{PA}}_{i}}$ as parents, formalizing the idea that the mechanisms $P_{X_{i}|\operatorname{{PA}}_{i}}$ also determine how the parents in $G$ influence the respective observation. Figure 1 shows $G^{m}$ for the DAG $X\to Y$. For our purpose, it is sufficient to consider $G^{1}$, where we have only one observation $x_{j}$ from each node $X_{j}$ in $G$. Accordingly, we construct $G^{1}$ from $G$ as follows: replace each variable $X_{j}$ with its observation $x_{j}$, and make $P_{X_{j}|\operatorname{{PA}}_{j}}$ a parent of $X_{j}$. Each $P_{X_{j}|\operatorname{{PA}}_{j}}$ becomes a root node, and these are the only root nodes in $G^{1}$. This placement of $P_{X_{j}|\operatorname{{PA}}_{j}}$ as root nodes (and thus algorithmically independent) formalizes the Independence of Mechanisms [^6] Janzing and Schölkopf [2010] call the algorithmic independence of different $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ Independence of Mechanisms. Here we prefer using this term for the underlying idea of representing them as root nodes.. Applying the algorithmic Markov condition to $G^{1}$ yields:
@@ 3 Decomposition of Randomness Deficiency
We start by presentation the decomposition of randomness deficiency in the bivariate case. Specifically, consider the DAG $X\rightarrow Y$ between a cause $X$ and its effect $Y$. Extending Eq. 4 with each distribution’s shortest representation, define the joint randomness deficiency of outcomes $(x,y)$ with respect to $P_{X,Y}$ by
@@
Lemma 3.1.
(Decomposition of randomness deficiency for a cause-effect pair) For any two random variables $X\rightarrow Y$ (i.e., $X$ being the cause of $Y$), and for specific observations $x$ and $y$, the following equality holds under Postulate 2.4 for the DAG in Figure 1:
@@
The proof relies on Lemma 2.5, and is provided in Section 10 of the supplementary.
@@
Figure 1: $x_{1},\dots,x_{m}$ sampled from $P_{X}$, $y_{1},\dots,y_{m}$ sampled from $P_{Y\mid X}$.
The following example shows why causal assumptions are needed for randomness deficiency to be additive.
Example 2 (No additivity without Independence of Mechanisms).
Let $\mathcal{X}=\mathcal{Y}=\{0,1\}^{d}$ and $P_{X}$ be the uniform distribution. Furthermore, let $P_{Y|X}$ be the deterministic mechanism
@@
where $\oplus$ denotes bitwise XOR and $x_{0}$ is an algorithmically random string with $K(x_{0})=d$. The mechanism $P_{Y\mid X}$ being deterministic implies that we always have $\delta(y\mid x)\stackrel{{\scriptstyle+}}{{=}}0$.
@@
We now generalize Lemma 3.1 to the multivariate case:
Theorem 3.2 (Decomposition of multivariate joint randomness deficiency).
Let the set of strings $x_{1},x_{2},\dots,x_{n}$ be causally connected by a directed acyclic graph $G$, so that the causal Markov condition holds for $G^{m}$. Then the joint randomness deficiency of all strings $x_{1},x_{2},\ldots,x_{n}$ decomposes into the conditional randomness deficiencies of the mechanisms:
@@
The proof is a simple induction over the number $n$ of nodes according to the causal ordering. The induction step from $n-1$ to $n$ uses Lemma 3.1 where $y:=x_{n}$ and $x$ is the single multivariate parent $x:=(x_{1},\dots,x_{n-1})$; see Section 10 in the supplementary for details.
4 Monotonicity of Randomness Deficiency
In particular, Theorem 3.2 allows us to draw the following straightforward conclusion:
Theorem 4.1 (Weak anomalies do not cause stronger ones).
@@ 4 Monotonicity of Randomness Deficiency
and the conditions of Theorem 3.2 are met,
[…]
The proof follows easily from Theorem 3.2 together with Corollary 4.1.11 from Gács [2021] which states non-increasingness of $\delta$ under marginalization and thus
$\delta(x_{i})\stackrel{{\scriptstyle+}}{{\leq}}\delta(x_{1},\dots,x_{n}).$
In essence, (7) states that none of the nodes displays a randomness deficiency that exceeds the conditional randomness deficiency of the root cause. In a nutshell, “weak outliers cannot cause strong ones”. – We have thus found an AIT version of Lemma 1.1.777Hence, in the language of modern resource theories in physics [Coecke et al., 2016], Definition 5.1., an anomaly is a “resource” and $\delta$ a “monotone” measuring its value. It is worth noting that Levin [1984] (see also Gács [2021]) demonstrated a similar principle, calling it randomness conservation. It says that the output of a mechanism cannot exhibit a substantially larger randomness deficiency than the input - under the condition that the mechanism itself has a constant description length. In physics, randomness deficiency corresponds to a lack of entropy [Zurek, 1989, Gács, 1994]. Therefore, the second law of thermodynamics amounts to an instance of randomness conservation [Ebtekar and Hutter, 2025].
In our context, simplicity of mechanisms is generalized by the independence of mechanisms. Example 2 illustrates why the independence of mechanisms principle is essential. It describes a scenario where a value $y$ with randomness deficiency $d$ emerges, yet this deficiency cannot be attributed to either the input $x$ or the mechanism generating $y$ from $x$.
@@ 5 Relation to computable anomaly scores
This section describes simple scenarios in which randomness deficiency boils down to simple and well-known scores. One may ask why one should start with an uncomputable concept in the first place to end up with something simple anyway. The answer is that the properties we have shown for randomness deficiency guide us in calibrating simple scores, with desirable properties such as decomposition of joint scores into scores of the mechanisms and monotonicity of scores.
@@
Let X be generated by a causal Bayesian network with linear structural equations $\textbf{X}=A\textbf{X}+{\bf N}$, where $A$ is strictly lower triangular and $N_{i}$ are independent noise variables with variance $\sigma_{i}^{2}$. Using transformation ${\bf N}=(I-A)\textbf{X}$, we obtain a diagonal form
@@
For large $n_{i}$, the expression $n_{i}^{2}/\sigma_{2}$ is roughly proportional to the randomness deficiency of the mechanism $P(X_{i}\mid PA_{i})$. Hence, decomposition of randomness deficiencies translates asymptotically into decompositions of the squared M-distance (as used for multivariate anomaly detection Aggarwal [2016]) into $z^{2}$-scores. Further, M-distance is non-increasing with respect to marginalization to a subset of variables (see Section 11 in the supplementary), resembling the monotonicity of randomness deficiency.
Example 4 supports Principle 1.2: a conservative bound on the randomness deficiency provides us with a computable e-value, as a function of the $z^{2}$-scores. Moreover, the conclusion that the $z^{2}$-score of any variable cannot exceed the sum of all these “noise scores” still holds, because our reasoning above is entirely based on linear algebra.
The following example shows that root causes may have smaller marginal randomness deficiency than their downstream effect:
@@
Let $X_{2}=2X_{1}+E_{2}$ with $X_{1},E_{2}\sim{\mathcal{N}}(0,1)$. Then $X_{2}\sim{\mathcal{N}}(0,\sqrt{3})$. Increasing $x_{2}$ by manual perturbation induces a marginal randomness deficiency that increases with $x_{2}^{2}/3$ for $x_{2}\to\infty$, while the conditional deficiency increases even like $x^{2}_{2}/1$. Thus, we have $\delta(x_{2}\mid x_{1})>\delta(x_{2})$ and $\delta(x_{1},x_{2})>\delta(x_{2})$ for large perturbations. If $X_{1}$ and $X_{2}$ together influence a third variable $X_{3}=X_{1}-X_{2}+E_{3}$ with $E_{3}\sim\mathcal{N}(0,1)$, we have $X_{3}\sim{\mathcal{N}}(0,\sqrt{2})$, thus $\delta(x_{3})$ increases with $x^{2}_{2}/2$. Hence, $\delta(x_{3})$ increases stronger than $\delta(x_{2})$, but weaker than $\delta(x_{2}\mid x_{1})$ and $\delta(x_{1},x_{2})$.
This paradox resembles Example 2.1 in Li et al. [2024], where it is phrased in terms of $z^{2}$-scores. Following Principle 1.2, we can describe it entirely in terms of M-distances (which are just $z^{2}$-scores in the one-dimensional case): the M-distance of $x_{3}$ can be larger than the M-distance of its root cause $x_{2}$, but not larger than the M-distance of the pair $(x_{1},x_{2})$, that is, the full cause of $x_{3}$. In other words, to fully quantify the anomaly introduced by perturbing $x_{2}$ we need to also account for the destroyed coupling between $x_{1}$ and $x_{2}$, not only for the size of the value $x_{2}$ itself.
The merit of algorithmic information theory is the insight that anomaly scores of the vector $(x_{1},x_{2})$ and the scalar $x_{3}$ are indeed comparable, and also comparable to the conditional score of $x_{2}$ given $x_{1}$, because their calibration was guided by the randomness deficiency.
@@
Here we can see that words with unexpectedly low or high Hamming weight, both are assigned high outlier scores, enabling us to compare both types of outliers. This works despite the probability mass function being monotonic in the Hamming weight $\ell$.
Multiple works [Aggarwal, 2016, Akoglu et al., 2012, Noble and Cook, 2003] propose detecting anomalies by compression, but consider anomalies that have higher compression length than usual, although randomness deficiency seems to flag only anomalies that have unexpectedly low compression length. Example 6 also resolves this paradox: For $p<1/2$, any string with $K(w)>m\cdot H_{2}(p)$ must have Hamming weight larger than $mp$. Accordingly, its randomness deficiency is bounded from below by (8). Hence, paradoxically, an unusually high compression length also implies non-zero randomness deficiency, because it entails an increase of the log likelihood term that outweighs the observed increase in compression length.
More sophisticated notions of anomalies can be obtained, for instance, when $X$ is graph-valued and an anomaly is given by large cliques [Aggarwal, 2016] – e.g. when fraudsters work together frequently on illegal activities, their communication is densely connected. To estimate the random deficiency of a graph $G$ with a clique of size $k$, one only needs to define a probability distribution on the set of graphs with $m$ nodes, and bound the description length of a graph by counting the number of graphs with cliques of size $k$.
@@ 6 Experiment with Lempel-Ziv Compression
So far we have accounted for the term $K(x|P^{*})$ only by very rough upper bounds rather than trying to approximate it. Here we describe a toy scenario in which we can detect anomalies using the Lempel-Ziv compression algorithm; more difficult scenarios may demand more powerful compression algorithms that are closer to general AI.
@@
7 Conclusion and Future Work
Algorithmic randomness deficiency offers a principled and flexible definition of outliers, without prior specification of the feature that exposes the anomaly. On a causal Bayesian network, we saw that the randomness deficiency of a joint observation decomposes along individual causal mechanisms, subject to the Independent Mechanisms Principle. This allows us to trace anomalous observations back to their root causes, identifying specific mechanisms that exhibit atypical behavior. Furthermore, we showed that weak outliers cannot be responsible for producing strong outliers, thus extending Levin’s law of randomness conservation. This foundational insight can help calibrating anomaly scores in a way that supports root cause analysis in complex systems.

### v3

@@ 2.4 Algorithmic Markov condition
Let $x_{1},\dots,x_{n}$ be binary words describing objects whose causal relations are given by the DAG $G$. Then, the joint complexity of $(x_{1},\dots,x_{n})$ decomposes as
@@
Equivalently, for any three sets $R,S,T$ of nodes such that $R$ d-separates $S$ and $T$, we have
[…]
Janzing and Schölkopf [2010] argue that, within a causal Bayesian network whose nodes represent random variables, the mechanisms $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ constitute independent information-bearing objects. Given $m$ observations of $n$-tuples $(x^{1}_{1},\dots,x_{n}^{1}),\dots,(x^{m}_{1},\dots,x_{n}^{m})$, they define a DAG $G^{m}$ on $n\times(m+1)$ nodes, corresponding to the $n$ mechanisms $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ and $n\times m$ observations $x_{j}^{l}$. The edges of $G^{m}$ consist of assigning the observations $\operatorname{{pa}}^{l}_{j}$ and the node $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ as parents of $x_{j}^{l}$. Figure 1 shows $G^{m}$, when $G$ is a cause-effect pair $X\to Y$.
For our purposes, it suffices to consider $G^{1}$, containing only one observation $x_{j}$ for each node $X_{j}$ in $G$. Accordingly, $G^{1}$ is constructed from $G$ as follows: replace each variable $X_{j}$ with its observation $x_{j}$, and add $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ as a parent of $X_{j}$. Each $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ becomes a root node, and these are the only root nodes in $G^{1}$.
When the mechanisms $P_{X_{j}\mid\operatorname{{PA}}_{j}}$ appear as root nodes, Postulate 2.4 implies their algorithmic independence from one another; Janzing and Schölkopf [2010] refer to this as Independence of Mechanisms. More generally, we can allow mechanisms to share information by adding a new root node as their shared parent. In either case, the algorithmic Markov condition can be interpreted as saying that multiple mechanisms cannot fail in a correlated manner. Instead, correlated anomalies must be attributable to a common root cause. To derive a formal statement, we apply Postulate 2.4 to $G^{1}$:
@@ 3 Decomposition of Randomness Deficiency
We start by presenting the decomposition of randomness deficiency in the bivariate case. Specifically, consider the DAG $X\to Y$ between a cause $X$ and its effect $Y$. Extending Eq. 4 to joint and conditional distributions, define the joint randomness deficiency of outcomes $(x,y)$ with respect to $P_{X,Y}$ by
@@
Lemma 3.1 (Decomposition of randomness deficiency for a cause-effect pair).
For any two random variables $X\to Y$ (i.e., $X$ being the cause of $Y$), and for individual observations $x$ and $y$, the following equality holds under Postulate 2.4 for the graph $G^{1}$ in Figure 1:
@@
The proof relies on Lemma 2.5, and is provided in Section 9 of the supplementary. The following example shows why the algorithmic Markov condition is essential in order for the randomness deficiency to be additive.
@@
Figure 1: The graph $G^{m}$, derived from $G$ of the form $X\to Y$, for $x^{1},\dots,x^{m}$ sampled from $P_{X}$ and $y^{1},\dots,y^{m}$ sampled from $P_{Y\mid X}$.
Example 2 (No additivity without the algorithmic Markov condition).
Let $\mathcal{X}=\mathcal{Y}=\{0,1\}^{d}$ and $P_{X}$ be the uniform distribution. Furthermore, let $P_{Y\mid X}$ be the deterministic mechanism
@@
where $\oplus$ denotes bitwise XOR and $x_{0}$ is an algorithmically random string with $K(x_{0})\stackrel{{\scriptstyle+}}{{=}}d$. Since $P_{Y\mid X}$ is deterministic, we always have $\delta(y\mid x)\stackrel{{\scriptstyle+}}{{=}}0$.
@@
The multivariate generalization of Lemma 3.1 is this paper’s main result:
Theorem 3.2 (Decomposition of joint randomness deficiency).
Let the set of strings $x_{1},x_{2},\dots,x_{n}$ be causally connected by a directed acyclic graph $G$, so that the algorithmic Markov condition holds for $G^{1}$. Then, the joint randomness deficiency of $x_{1},x_{2},\ldots,x_{n}$ decomposes into a sum of conditional randomness deficiencies of the mechanisms:
@@
The proof is by induction over the number $n$ of nodes; in order to chain the $\stackrel{{\scriptstyle+}}{{=}}$ equalities, we treat $n$ as a constant. See Section 9 in the supplementary for details, including a proof of the following corollary:
Corollary 3.3 (Weak outliers do not cause stronger ones).
@@
and the conditions of Theorem 3.2 are met, then
$\delta(x_{i})\stackrel{{\scriptstyle+}}{{\leq}}\delta(x_{j}\mid\operatorname{{pa}}_{j})\quad\forall i\in\{1,\dots,n\}.$ (7)
On the bivariate graph $X\to Y$ whose unique root anomaly is $X$, Corollary 3.3 implies $\delta(y)\stackrel{{\scriptstyle+}}{{\leq}}\delta(x)$. More generally, (7) states that no node has a randomness deficiency exceeding that of the root cause. In a nutshell, “weak outliers cannot cause strong ones”. – We have thus found an AIT version of Lemma 1.1.666Hence, in the language of modern resource theories in physics ([Coecke et al., 2016], Definition 5.1), an anomaly is a “resource”, and $\delta$ is a “monotone” measuring its value.
It is worth noting that Levin [1984] (see also Gács [2021]) demonstrated a similar principle, calling it randomness conservation. It says that a mechanism’s typical output cannot exhibit a substantially larger randomness deficiency than its input - under the condition that the mechanism is simple, i.e., has constant description length. In physics, randomness deficiency corresponds to a lack of entropy [Zurek, 1989, Gács, 1994]. Therefore, the second law of thermodynamics amounts to an instance of randomness conservation [Ebtekar and Hutter, 2025]. By conditioning the randomness deficiencies on the mechanisms, we obtained results that hold even when the mechanisms are not simple.
@@ 4 Relation to computable anomaly scores
In practice, we cannot directly use the randomness deficiency to detect outliers, because it is not computable. However, it is semicomputable: by searching over descriptions of $x$, one can gradually improve the upper bound on its Kolmogorov complexity, thus obtaining approximations that monotonically approach the randomness deficiency in the limit. Since a generic search would be extremely slow, the search should focus on features or models suitable to the domain in question. We leave the development of good domain-specific anomaly search algorithms to future work.
Alternatively, the search can be performed manually without a computer algorithm. During statistical analysis, one can identify any computable feature, perhaps even after observing $x$, and from it derive a lower bound on the randomness deficiency.
This section describes simple scenarios in which such bounds boil down to simple and well-known outlier scores. Thus, the randomness deficiency provides a practical recipe for obtaining and calibrating computable scores. Moreover, its theoretical properties motivate us to seek computable scores with the same desirable properties: decomposition of joint scores into scores of the mechanisms, and monotonicity under marginalization.
@@
Let X be generated by a causal Bayesian network with linear structural equations $\textbf{X}=A\textbf{X}+{\bf N}$, where $A$ is strictly lower triangular and $N_{i}$ are independent noise variables with variance $\sigma_{i}^{2}$. Using the transformation ${\bf N}=(I-A)\textbf{X}$, we obtain a diagonal form
@@
For large $n_{i}$, the expression $n_{i}^{2}/\sigma_{2}$ is roughly proportional to the randomness deficiency of the mechanism $P(X_{i}\mid PA_{i})$. Hence, decomposition of randomness deficiencies translates asymptotically to decompositions of the squared M-distance (as used for multivariate anomaly detection Aggarwal [2016]) into $z^{2}$-scores. Furthermore, M-distance is non-increasing with respect to marginalization to a subset of variables (see Section 10 in the supplementary), resembling the monotonicity of randomness deficiency.
Example 4 supports Principle 1.2: a conservative bound on the randomness deficiency yields a computable e-value, as a function of the $z^{2}$-scores. Moreover, we verified using linear algebra that the $z^{2}$-score of any variable cannot exceed the sum of all the “noise scores”.
Note that the upper bound in Corollary 3.3 is a conditional randomness deficiency. The following example shows that a root cause may have a smaller marginal randomness deficiency than its downstream effects.
@@
With $N_{i}\sim{\mathcal{N}}(0,1)$, consider the three-node model
$X_{1}$ $:=N_{1},$
$X_{2}$ $:=2X_{1}+N_{2},$
$X_{3}$ $:=X_{1}-X_{2}+N_{3}.$
Then, $X_{2}\sim{\mathcal{N}}(0,\sqrt{5})$. Suppose $n_{1},n_{3}$ take on typical values, while $n_{2}$ is anomalously large. Then, the marginal randomness deficiency $\delta(x_{2})$ increases with $n_{2}^{2}/5$, while the conditional randomness deficiency $\delta(x_{2}\mid x_{1})$ increases with $n^{2}_{2}/1$ (where we’ve ignored the factor of $\frac{\log e}{2}$ from Example 3). Since $X_{3}\sim{\mathcal{N}}(0,\sqrt{3})$, $\delta(x_{3})$ increases with $n^{2}_{2}/3$. Thus, for large perturbations of $x_{2}$, we have $\delta(x_{2})\ll\delta(x_{3})\ll\delta(x_{2}\mid x_{1})\approx\delta(x_{1},x_{2})$. Looking only at marginal scores, it would seem as though $x_{2}$ caused a stronger outlier $x_{3}$. Nonetheless, $\delta(x_{3})\stackrel{{\scriptstyle+}}{{\leq}}\delta(x_{2}\mid x_{1})$ in agreement with Corollary 3.3.
This paradox resembles Example 2.1 in Li et al. [2024], where it is phrased in terms of $z^{2}$-scores. Following Principle 1.2, we can describe it entirely in terms of M-distances (which are just $z^{2}$-scores in the one-dimensional case): the M-distance of $x_{3}$ can be larger than the M-distance of its root cause $x_{2}$, but not larger than the M-distance of the pair $(x_{1},x_{2})$, i.e., the full cause of $x_{3}$. In other words, to fully quantify the anomaly introduced by perturbing $x_{2}$, we must also account for the destroyed coupling between $x_{1}$ and $x_{2}$, not only for the size of the value $x_{2}$ itself.
The insight from algorithmic information theory is that the anomaly scores of the vector $(x_{1},x_{2})$ and the scalar $x_{3}$ are indeed comparable, and also comparable to the conditional score of $x_{2}$ given $x_{1}$, because their calibration was guided by the randomness deficiency.
@@
Here we see that words with unexpectedly low or unexpectedly Hamming weight are both assigned high outlier scores, enabling us to compare both types of outliers. This works despite the probability mass function being monotonic in the Hamming weight $\ell$ [^7] Cover and Thomas [2006], in Section 11.2, call a word “$\epsilon$-typical” when it satisfies ${\operatorname{KL}\left({\frac{\ell}{m}\mid\mid p}\right)}\leq\epsilon$. Their equation (11.67) finds that such a word occurs with probability at least $1-(m+1)^{2}\cdot e^{-\epsilon m}$.
Many works [Aggarwal, 2016, Akoglu et al., 2012, Noble and Cook, 2003] propose detecting anomalies by compression, but consider anomalies that have higher compression length than usual, seemingly in conflict with the randomness deficiency which flags low compression lengths. Example 6 resolves this paradox: for $p<1/2$, any word with $K(w)>m\cdot H(p)$ must have Hamming weight larger than $mp$. Accordingly, its randomness deficiency is bounded from below by (8). Hence, paradoxically, an unusually high compression length also implies non-zero randomness deficiency, because it entails an increase of the log likelihood term that outweighs the observed increase in compression length.
More sophisticated notions of anomalies can be obtained, for instance, when $X$ is graph-valued and an anomaly is given by large cliques [Aggarwal, 2016] – e.g. when fraudsters work together frequently on illegal activities, their communication is densely connected. To estimate the random deficiency of a graph $G$ with a clique of size $k$, it suffices to define a probability distribution on the set of graphs with $m$ nodes, and bound a graph’s description length by counting the number of graphs with cliques of size $k$.
@@ 5 Experiment with Lempel-Ziv Compression
So far we have accounted for the term $K(x\mid P^{*})$ only by very rough upper bounds rather than trying to approximate it. Here we describe a toy scenario in which we can detect anomalies using the Lempel-Ziv compression algorithm; more difficult scenarios may demand more powerful compression algorithms that are closer to general AI.
@@
6 Conclusions
The algorithmic randomness deficiency offers a principled unified definition of outliers, without prior specification of the feature that exposes the anomaly. On a causal Bayesian network, we saw that the randomness deficiency of a joint observation decomposes along individual causal mechanisms, subject to the Independent Mechanisms Principle. This allows us to trace anomalous observations back to their root causes, identifying specific mechanisms responsible for the anomaly. Moreover, we showed that weak outliers cannot cause stronger outliers, extending Levin’s law of randomness conservation. These foundational insights can help calibrate anomaly scores as well as inspire more versatile anomaly detection algorithms, supporting root cause analysis in complex systems.